Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
Answer by stanbon(75887)

Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
-------------------------
Upstream DATA:
distance = 5 miles ; rate = b-4 mph ; time = d/r = 5/(b-4) hrs.
----------------
Downstream DATA:
distance = 5 miles ; rate = b+4 mph ; time = 5/(b+4) hrs.
-----------------
Equation:
up - down = (1/5) hr
5/(b-4) - 5/(b+4) = 1/5
-----
25(b+4) - 25(b-4) = (b^2-16)
100+100 = b^2-16
b^2-216 = 0
Positive solution:
b = sqrt(216)
b = 14.7 mph (speed of the boat in quiet water)
===================
Cheers,
Stan H.